# InvZW: Invariant Feature Learning via Noise-Adversarial Training for Robust Image Zero-Watermarking

### v1

@@ Abstract
Index Terms:
Robust Zero-watermarking, invariant feature learning, adversarial training, representation learning, deep watermarking
I Introduction
Image watermarking is a widely used technique for embedding hidden information into digital media for purposes such as copyright protection and content authentication. In contrast to standard watermarking methods that modify the cover image by embedding a watermark directly into its pixel or transform domain representation [1], zero-watermarking leaves the original image unchanged. Instead, it extracts robust features from the image and associates them with a watermark to form a signature, which is stored externally without altering the image. During usage, features extracted from a potentially marked image are used to either verify the watermark or recover it. Zero-watermarking is particularly useful in domains where modifying the original image is undesirable, such as in medical diagnostics, legal evidence, and satellite imaging. There are two primary objectives of zero-watermarking: verification-based approaches, which confirm the presence of a watermark [2], and multibit extraction-based methods, which aim to retrieve the full watermark content from the extracted features [3].
@@ II Related Work
II-A Invariant Feature Learning
@@
II-B Deep Learning-based Image Zero-Watermarking
@@
Beyond zero-watermarking, several deep image watermarking schemes [7, 8] have explored the use of pre-trained networks—particularly self-supervised models—as fixed feature extractors. These methods embed watermarks by optimizing the input image while keeping the model weights frozen, allowing the watermark to be revealed in the feature space. By leveraging the inherent robustness of self-supervised representations, these approaches achieve resilience to common distortions without the need for retraining the feature extractor.
@@ III Proposed Method
III-A Module 1: Invariant Feature Learning via Noise-Adversarial Training
Fig. 3 illustrates the proposed invariant feature learning module based on noise-adversarial training. The framework consists of three main components: a feature extractor, a discriminator, and a reconstructor. Given both original and distorted images, the feature extractor—implemented using a Transformer-based architecture—generates high-level representations intended to be invariant to distortions. The discriminator is trained to classify whether a given feature originates from an original or distorted image, while the feature extractor is trained adversarially to fool the discriminator. This minimax interaction encourages the extractor to produce features that are indistinguishable across distortions, thereby promoting invariance. To ensure that the learned features retain semantic content, the reconstructor is trained to reconstruct the original image from these features, forcing the invariant features to preserve meaningful information relevant to the image itself.
@@ III-A1 Feature Extractor
From the final Transformer block, we extract the output corresponding to the [CLS] token as the global representation of the image, denoted as $F=\mathrm{FE}(x)=z^{(L)}_{\text{CLS}}$. Both the original and distorted images are passed through the same feature extractor, and the model will be trained using adversarial objectives to ensure that the extracted features remain invariant to distortions. These features are then passed to the discriminator and reconstructor for further supervision, as detailed in the following subsections.
III-A2 Discriminator
@@
Given an input image of size $128\times 128\times 3$, the discriminator first divides it into 64 non-overlapping patches of size $16\times 16$. Each patch is flattened and linearly projected into a 64-dimensional latent space using a learnable embedding layer. To preserve spatial ordering, positional embeddings are added to each projected patch vector. A learnable [CLS] token is prepended to the patch sequence, serving as a global aggregation token for classification.
The sequence, including the [CLS] token, is passed through a stack of $L=8$ Transformer blocks. Each block contains a MHSA layer followed by a FFN as in Fig. 3, with residual connections and layer normalization applied at each submodule. This architecture allows the model to capture both local and global relationships across the image patches.
After Transformer processing, the output corresponding to the [CLS] token is extracted and passed through a dense classification head consisting of a single linear layer followed by a sigmoid activation: $D(F)=\sigma(W_{d}\cdot F_{\text{CLS}}+b_{d}),$ where $F_{\text{CLS}}$ is the final embedding of the classification token, $W_{d}\in\mathbb{R}^{1\times d}$ and $b_{d}\in\mathbb{R}$ are the learnable weight and bias parameters of the classification head, and $\sigma$ denotes the sigmoid function.
The discriminator will be trained to distinguish between features derived from original and distorted images. Its gradients provide adversarial feedback to the feature extractor, encouraging the extractor to produce representations that are invariant with respect to the applied distortions.
III-A3 Reconstructor
The reconstructor, illustrated in Fig. 4, can be conceptually described as an autoencoder-like architecture designed to enforce semantic preservation in the learned features. It consists of an encoder path, a bottleneck, and a decoder path. The input to the reconstructor is the feature tensor $F\in\mathbb{R}^{128\times 128\times 3}$ produced by the feature extractor. This reconstruction objective ensures that the extracted features retain sufficient semantic information, discouraging the feature extractor from collapsing into trivial or non-informative representations.
@@ III-A4 Training
0:  Feature Extractor $\mathrm{FE}$, Discriminator $D$, Reconstructor $R$, transformation function $T$, dataset $X$, batch size $B$, learning rate $\alpha$, loss weights $\lambda_{1},\lambda_{2}$.
1:  for each epoch do
2:     for each batch of images $x$ sampled from $X$ do
3:        Generate distorted images: $x^{\prime}=T(x)$.
4:        Extract features: $F=\mathrm{FE}(x)$, $F^{\prime}=\mathrm{FE}(x^{\prime})$.
5:        Train Discriminator:
6:        Compute labels: $y=1$ for $F$, $y=0$ for $F^{\prime}$.
7:        Compute discriminator loss $\mathcal{L}_{D}$.
8:        Update $D$ using $\theta_{D}\leftarrow\theta_{D}-\alpha\nabla_{\theta_{D}}\mathcal{L}_{D}$.
9:        Train Feature Extractor:
10:        Compute adversarial loss $\mathcal{L}_{\mathrm{adv}}$.
11:        Compute reconstruction loss $\mathcal{L}_{R}$.
12:        Compute total loss $\mathcal{L}_{\mathrm{FE}}=\lambda_{1}\mathcal{L}_{\mathrm{adv}}+\lambda_{2}\mathcal{L}_{R}$.
13:        Update $\mathrm{FE}$ using $\theta_{\mathrm{FE}}\leftarrow\theta_{\mathrm{FE}}-\alpha\nabla_{\theta_{\mathrm{FE}}}\mathcal{L}_{\mathrm{FE}}$.
14:        Train Reconstructor:
15:        Reconstruct images: $\hat{x}=R(F)$.
16:        Compute reconstruction loss $\mathcal{L}_{R}$.
17:        Update $R$ using $\theta_{R}\leftarrow\theta_{R}-\alpha\nabla_{\theta_{R}}\mathcal{L}_{R}$.
18:     end for
19:     Save model weights every 10 epochs.
20:  end for
III-A5 Theoretical Justification
The proposed noise-adversarial training framework is theoretically supported by two complementary perspectives: the Information Bottleneck (IB) principle [23] and domain adversarial learning [24].
First, we interpret the dual-objective formulation—adversarial training for invariance and reconstruction for semantic preservation—through the lens of the Information Bottleneck principle. The goal is to learn a feature representation $F$ that maximizes mutual information with the input image $x$, while minimizing mutual information with irrelevant nuisance factors introduced by distortion $x^{\prime}$. This trade-off can be formalized as:
@@ III-A5 Theoretical Justification
where $I(\cdot,\cdot)$ denotes mutual information, and $\beta>0$ balances informativeness against invariance. While mutual information is not computed explicitly, the adversarial loss serves as a surrogate to minimize $I(F,x^{\prime})$ by aligning the feature distributions of $x$ and $x^{\prime}$, i.e., $p(F)\approx p(F^{\prime})$. Simultaneously, the reconstruction loss enforces that $F$ retains sufficient information to recover the original content, implicitly maximizing $I(F,x)$.
This framework can also be viewed as a case of domain adversarial learning, where the distortion acts as a latent domain shift. In this view, the original image $x$ and its distorted version $x^{\prime}$ belong to two domains with shared semantic content but domain-specific variations (i.e., noise). The adversarial training objective ensures that the learned features are indistinguishable across domains:
$\mathbb{E}_{x\sim p(x)}\left[D(\mathrm{FE}(x))\right]=\mathbb{E}_{x^{\prime}\sim p_{T}(x)}\left[D(\mathrm{FE}(x^{\prime}))\right]=0.5.$ (6)
This alignment encourages the learned representation $F$ to be domain-invariant (i.e., distortion-invariant), such that noise does not affect the feature encoding. However, domain invariance alone is insufficient, as it may lead to degenerate solutions where the extracted features collapse to trivial or constant vectors. To prevent such feature collapse, the reconstructor imposes a functional constraint: $x\approx R(\mathrm{FE}(x)),\quad x^{\prime}\approx R(\mathrm{FE}(x^{\prime})),$ which enforces that the features remain semantically meaningful and non-trivial. Together, these constraints ensure that the feature extractor learns robust, invariant, and expressive representations—properties that are critical for downstream watermark retrieval in zero-watermarking applications.
III-B Module 2: Learning-based Zero-Watermarking with Invariant Feature
@@
III-B1 Feature Extraction and Encoding
@@
III-B2 Loss Function and Optimization
@@ III-B3 Watermark Extraction
III-B4 Method Advantages
A key practical advantage of the proposed formulation is its flexibility in associating different watermark messages with the same image. Since the reference signature $C$ is optimized in feature space and not embedded in the image, it can be re-trained per user or application without modifying the image or the invariant feature extractor. This supports a one-to-many mapping: a single image can be linked to multiple watermark identities, each identified by a unique $C$. During extraction, only the correct reference signature will yield a successful extraction, providing a mechanism for ownership differentiation, user-specific claims, or multi-level authorization.
From a theoretical perspective, the proposed approach can be interpreted as solving a constrained optimization problem in the invariant feature space. Given a fixed, distortion-invariant representation $\tilde{F}$, the training of $C$ seeks to project $\tilde{F}$ onto a set of learned directional codes $\{C_{1},\dots,C_{k}\}$ such that their inner products reflect a target binary vector $W$. The sigmoid activation and binary cross-entropy loss induce a soft margin between watermark bits, effectively identifying a hyperplane that separates $\tilde{F}\cdot C_{i}\geq 0.5$ from $<0.5$. By minimizing this loss, the solution converges to a directionally optimal configuration of $C$, tailored to the unique geometry of the input feature. This learning-based formulation generalizes well to unseen transformations, as the optimization finds reference codes that are not just memorized, but robustly aligned with the semantic content captured by $\tilde{F}$.
@@ III-B4 Method Advantages
1:  Input: Original images $\{I_{i}\}_{i=1}^{N}$, Watermarks $\{W_{i}\}_{i=1}^{N}$, Reference $C\in\mathbb{R}^{k\times d}$, Feature Extractor $FE$
2:  Output: Optimized Reference $C\in\mathbb{R}^{k\times d}$ and $\Psi$
@@
4:     Extract invariant feature map: $F=FE(I_{i})\in\mathbb{R}^{128\times 128\times 3}$
5:     Compute $\tilde{F}=\Psi(F),\quad\tilde{F}\in\mathbb{R}^{d}$
6:     for each bit $b=1$ to $B$ do
7:        Compute dot product: $s_{i,b}=\tilde{F}_{i}\cdot C[i,b,:]$
@@
10:     Compute Binary Cross Entropy loss: $\mathcal{L}_{W}=\text{BCE}(\hat{W}_{i},W_{i})$
@@
12:     Total Loss: $\mathcal{L}_{total}$= $\mathcal{L}_{W}+\mathcal{L}_{C}$
13:     Update $C$ using $\theta_{C}\leftarrow\theta_{C}-\alpha\nabla_{\theta_{C}}\mathcal{L}_{total}$.
14:     Update $\Psi$ using $\theta_{\Psi}\leftarrow\theta_{\Psi}-\alpha\nabla_{\theta_{\Psi}}\mathcal{L}_{total}$.
@@
16:  Watermark Extraction:
17:  Input: Noisy image $I^{\prime}$, Feature Extractor $FE$, Optimized Reference $C$, Optimized $\Psi$
18:  Extract invariant feature map: $F=FE(I_{i})$
19:  Apply $\Psi$: $\tilde{F}=\Psi(F)$
20:  for each bit $b=1$ to $B$ do
21:     Compute dot product: $s^{\prime}_{b}=\tilde{F}\cdot C[*,b,:]$
22:     Apply sigmoid: $\hat{W}_{b}=\sigma(s^{\prime}_{b})$
23:     Assign bit: $W_{b}=\begin{cases}1&\text{if }\hat{w}_{b}\geq 0.5\\ 0&\text{otherwise}\end{cases}$
24:  end for
25:  Return extracted watermark $W$
IV Experimental Result
This section presents evaluation of the proposed framework. Section IV-A describes the datasets and preprocessing and augmentation strategies used for training and evaluation. Section IV-B analyzes the quality and stability of the learned invariant features, including feature consistency under distortion and the quality of image reconstruction. Section IV-C compares our method with existing self-supervised feature extractors, such as SimCLR, BYOL, and VICReg, using cosine similarity and linear evaluation metrics to assess feature robustness. Finally, Section IV-D demonstrates the effectiveness of our learning-based zero-watermarking scheme, providing comparisons with competitive baselines under a range of distortions.
IV-A Dataset and Feature Training

### v2

@@ Abstract
Keywords Deep robust zero-watermarking $\cdot$ invariant feature learning $\cdot$ adversarial training $\cdot$ representation learning
1 Introduction
Image watermarking is a widely used technique for embedding hidden information into digital media for purposes such as copyright protection and content authentication [29]. In contrast to standard watermarking methods that modify the cover image by embedding a watermark directly into its pixel or transform domain representation, zero-watermarking leaves the original image unchanged. Instead, it extracts robust features from the image and associates them with a watermark to form a signature, which is stored externally without altering the image. During usage, features extracted from a potentially marked image are used to either verify the watermark or recover it. This approach ensures that the visual integrity of the original image remains intact, making it ideal for scenarios where any modification is impermissible. While zero-watermarking has traditionally been associated with highly sensitive use cases, its non-intrusive and robust nature makes it broadly applicable. By enabling reliable copyright protection and content verification without altering the original content, it offers a practical and versatile solution across a wide range of image-based applications. There are two primary objectives of zero-watermarking: verification-based approaches, which confirm the presence of a watermark [18], and multibit extraction-based methods, which aim to retrieve the full watermark content from the extracted features [8].
@@ 2 Related Work
2.1 Invariant Feature Learning
@@
2.2 Deep Learning-based Image Zero-Watermarking
@@
Beyond zero-watermarking, several deep image watermarking schemes [24, 7] have explored the use of pre-trained networks, particularly self-supervised models—as fixed feature extractors. These methods embed watermarks by optimizing the input image while keeping the model weights frozen, allowing the watermark to be revealed in the feature space. By leveraging the inherent robustness of self-supervised representations, these approaches achieve resilience to common distortions without the need for retraining the feature extractor.
@@ 3 Proposed Method
3.1 Module 1: Invariant Feature Learning via Noise-Adversarial Training
Fig. 3 illustrates the proposed invariant feature learning module based on noise-adversarial training. The framework consists of three main components: a feature extractor, a discriminator, and a reconstructor. Given both original and distorted images, the feature extractor, implemented using a Transformer-based architecture, generates high-level representations intended to be invariant to distortions. The discriminator is trained to classify whether a given feature originates from an original or distorted image, while the feature extractor is trained adversarially to fool the discriminator. This minimax interaction encourages the extractor to produce features that are indistinguishable across distortions, thereby promoting invariance. To ensure that the learned features retain semantic content, the reconstructor is trained to reconstruct the original image from these features, forcing the invariant features to preserve meaningful information relevant to the image itself.
@@ 3.1.1 Feature Extractor
The output of the final Transformer block is a sequence of encoded patch tokens $\{z_{1}^{(L)},z_{2}^{(L)},\dots,z_{N}^{(L)}\}$, which collectively represent the image in a latent feature space. Both the original and distorted images are passed through the same feature extractor, and the model is trained using adversarial objectives to ensure that these learned representations remain invariant to distortions.
To enable spatial supervision, the sequence of encoded patch tokens is further linearly projected and reshaped into a spatial feature tensor $F_{s}\in\mathbb{R}^{128\times 128\times 3}$. This transformation is implemented via a linear projection applied to each patch token, followed by reshaping the sequence into a 2D grid aligned with the input resolution. The resulting spatial feature tensor is derived from the learned transformer representations (rather than the raw image) and is used as input to the discriminator and reconstructor modules described in the following subsections.
3.1.2 Discriminator
@@
Given the spatial feature tensor $F_{s}\in\mathbb{R}^{128\times 128\times 3}$ derived from encoded patch features, the discriminator first divides it into 64 non-overlapping patches of size $16\times 16$. Each patch is flattened and linearly projected into a 64-dimensional latent space using a learnable embedding layer. To preserve spatial ordering, positional embeddings are added to each projected patch vector.
The sequence of patch embeddings is passed through a stack of $L=8$ Transformer blocks. Each block contains a MHSA layer followed by a FFN as in Fig. 3, with residual connections and layer normalization applied at each submodule. This architecture allows the model to capture both local and global relationships across the image patches.
After Transformer processing, the output corresponding to the classification token is extracted and passed through a dense classification head with a softmax activation: $D(F)=\mathrm{Softmax}(W_{d}\cdot F_{\text{CLS}}+b_{d}),$ where $F_{\text{CLS}}$ denotes the final embedding of the classification token, and $W_{d}\in\mathbb{R}^{1\times d}$ and $b_{d}\in\mathbb{R}$ are learnable parameters.
The discriminator is trained to distinguish between features derived from original and distorted images. Its gradients provide adversarial feedback to the feature extractor, encouraging the extractor to produce representations that are invariant to the applied distortions.
3.1.3 Reconstructor
To prevent feature collapse, where the feature extractor outputs a constant or trivial representation (e.g., all zeros) regardless of the input, we introduce a reconstructor module $R$ as an auxiliary constraint. While such representations may appear invariant, they are semantically meaningless and unusable for watermarking. By enforcing that the extracted features must support accurate reconstruction of the original image, we ensure that the representations retain semantic information in addition to distortion invariance. If the features collapse, reconstruction fails with high error, which the model is penalized for, thereby effectively discouraging degenerate solutions. The reconstructor, illustrated in Fig. 4, can be conceptually described as an autoencoder-like architecture designed to enforce semantic preservation in the learned features. It consists of an encoder path, a bottleneck, and a decoder path. The input to the reconstructor is the feature tensor $F\in\mathbb{R}^{128\times 128\times 3}$ produced by the feature extractor.
@@ 3.1.4 Training
1:Feature Extractor $\mathrm{FE}$, Discriminator $D$, Reconstructor $R$, transformation function $T$, dataset $X$, batch size $B$, learning rate $\alpha$, loss weights $\lambda_{1},\lambda_{2}$
2:for each epoch do
3:  for each batch of images $x$ sampled from $X$ do
4:   Generate distorted images: $x^{\prime}=T(x)$
[…]
6:   Train Discriminator:
7:   Compute labels: $y=1$ for $F$, $y=0$ for $F^{\prime}$
8:   Compute discriminator loss $\mathcal{L}_{D}$
9:   Update $D$ using $\theta_{D}\leftarrow\theta_{D}-\alpha\nabla_{\theta_{D}}\mathcal{L}_{D}$
10:   Train Feature Extractor:
11:   Compute adversarial loss $\mathcal{L}_{\mathrm{adv}}$
12:   Compute reconstruction loss $\mathcal{L}_{R}$
13:   Compute total loss $\mathcal{L}_{\mathrm{FE}}=\lambda_{1}\mathcal{L}_{\mathrm{adv}}+\lambda_{2}\mathcal{L}_{R}$
14:   Update $\mathrm{FE}$ using $\theta_{\mathrm{FE}}\leftarrow\theta_{\mathrm{FE}}-\alpha\nabla_{\theta_{\mathrm{FE}}}\mathcal{L}_{\mathrm{FE}}$
15:   Train Reconstructor:
16:   Reconstruct images: $\hat{x}=R(F)$
17:   Compute reconstruction loss $\mathcal{L}_{R}$
18:   Update $R$ using $\theta_{R}\leftarrow\theta_{R}-\alpha\nabla_{\theta_{R}}\mathcal{L}_{R}$
19:  end for
20:  Save model weights every 10 epochs
21:end for
3.1.5 Theoretical Perspective
The proposed noise-adversarial training framework can be interpreted through two complementary perspectives: the Information Bottleneck (IB) principle [23] and domain adversarial learning [9].
From the perspective of the Information Bottleneck, the learning objective can be viewed as encouraging a representation $F$ that retains information about the underlying image content while reducing sensitivity to distortion. This can be conceptually related to the trade-off:
@@ 3.1.5 Theoretical Perspective
where $x$ denotes the original image and $x^{\prime}$ its distorted version. In our framework, this objective is not optimized explicitly. Instead, the adversarial loss encourages invariance by promoting similarity between feature distributions of $x$ and $x^{\prime}$, while the reconstruction loss encourages the representation to preserve semantically relevant information.
From a domain adversarial learning perspective, the distortion process can be interpreted as inducing a domain shift. The original image $x$ and its distorted version $x^{\prime}$ may be viewed as samples from related domains that share semantic structure but differ in nuisance factors. The adversarial objective encourages the feature extractor to produce representations that are difficult to distinguish across these domains. This can be interpreted as promoting approximate alignment between feature distributions:
$\mathbb{E}_{x\sim p(x)}\left[D(\mathrm{FE}(x))\right]\approx\mathbb{E}_{x^{\prime}\sim p_{T}(x)}\left[D(\mathrm{FE}(x^{\prime}))\right].$ (6)
However, enforcing invariance alone may lead to degenerate representations. The reconstruction constraint serves as a complementary mechanism that encourages the learned features to remain informative and structured:
$x\approx R(\mathrm{FE}(x)),\quad x^{\prime}\approx R(\mathrm{FE}(x^{\prime})),$ (7)
thereby discouraging trivial or collapsed solutions. Taken together, these perspectives provide conceptual motivation for the proposed design, suggesting that the combination of adversarial alignment and reconstruction encourages the learning of invariant yet expressive feature representations for robust watermarking.
3.2 Module 2: Learning-based Zero-Watermarking with Invariant Feature
@@
3.2.1 Feature Extraction and Encoding
@@
3.2.2 Loss Function and Optimization
@@ 3.2.3 Watermark Extraction
1:Original images $\{I_{i}\}_{i=1}^{N}$, watermarks $\{W_{i}\}_{i=1}^{N}$, reference tensor $C\in\mathbb{R}^{N\times B\times d}$, feature extractor $\mathrm{FE}$
2:Optimized reference tensor $C$ and projection module $\Psi$
@@
4:  Extract invariant feature map: $F_{i}=\mathrm{FE}(I_{i})$
5:  Compute projected feature: $\tilde{F}_{i}=\Psi(F_{i}),\ \tilde{F}_{i}\in\mathbb{R}^{d}$
6:  for $b=1$ to $B$ do
7:   Compute dot product: $s_{i,b}=\tilde{F}_{i}^{\top}C_{i,b}$
@@
10:  Compute BCE loss: $\mathcal{L}_{W}=\mathrm{BCE}(\hat{W}_{i},W_{i})$
@@
12:  Total loss: $\mathcal{L}_{\text{total}}=\mathcal{L}_{W}+\mathcal{L}_{C}$
13:  Update $C$: $\theta_{C}\leftarrow\theta_{C}-\alpha\nabla_{\theta_{C}}\mathcal{L}_{\text{total}}$
14:  Update $\Psi$: $\theta_{\Psi}\leftarrow\theta_{\Psi}-\alpha\nabla_{\theta_{\Psi}}\mathcal{L}_{\text{total}}$
@@
16:
17:Watermark Extraction
18:Input: Noisy image $I^{\prime}$, $\mathrm{FE}$, optimized $C$, optimized $\Psi$
19:Extract feature: $F^{\prime}=\mathrm{FE}(I^{\prime})$
20:Project: $\tilde{F}^{\prime}=\Psi(F^{\prime})$
21:for $b=1$ to $B$ do
22:  Compute: $s^{\prime}_{b}=(\tilde{F}^{\prime})^{\top}C_{b}$
23:  Apply sigmoid: $\hat{W}_{b}=\sigma(s^{\prime}_{b})$
24:  if $\hat{W}_{b}\geq 0.5$ then
25:   $W_{b}\leftarrow 1$
26:  else
27:   $W_{b}\leftarrow 0$
28:  end if
29:end for
30:return $W$
3.2.4 Method Advantages
A key practical advantage of the proposed framework is its flexibility in associating multiple watermark messages with a single image. Since the reference signature C is optimized in the invariant feature space rather than embedded in the image, it can be independently learned for each pair of image-watermarks without modifying the original image or retraining the feature extractor. In our formulation, a reference signature tensor of size $128\times 128\times 3$ is stored for each such pair. This enables a natural one-to-many mapping, where a single image can be linked to multiple watermark identities, each defined by a distinct reference signature. The storage requirement scales linearly with the number of assigned watermarks. During verification, the feature extracted from a query image is compared against candidate reference signatures, and only the correct signature produces a successful match, allowing reliable watermark identification and supporting user-specific ownership or multi-level authorization.
From a theoretical perspective, the proposed approach can be interpreted as solving a constrained optimization problem in the invariant feature space. Given a fixed, distortion-invariant representation $\tilde{F}$, the training of $C$ seeks to project $\tilde{F}$ onto a set of learned directional codes $\{C_{1},\dots,C_{k}\}$ such that their inner products reflect a target binary vector $W$. The sigmoid activation and binary cross-entropy loss induce a soft margin between watermark bits, effectively identifying a hyperplane that separates $\tilde{F}\cdot C_{i}\geq 0.5$ from <0.5<0.5. By minimizing this loss, the solution converges to a directionally optimal configuration of $C$, tailored to the unique geometry of the input feature. This learning-based formulation generalizes well to unseen transformations, as the optimization finds reference codes that are not just memorized, but robustly aligned with the semantic content captured by $\tilde{F}$.
4 Experimental Result
This section presents evaluation of the proposed framework. Section 4.1 describes the datasets and preprocessing and augmentation strategies used for training and evaluation. Section 4.2 analyzes the quality and stability of the learned invariant features, including feature consistency under distortion and the quality of image reconstruction. Section 4.3 compares our method with existing self-supervised feature extractors, such as SimCLR, BYOL, and VICReg, using cosine similarity and linear evaluation metrics to assess feature robustness. Finally, Section 4.4 demonstrates the effectiveness of our learning-based zero-watermarking scheme, providing comparisons with competitive baselines under a range of distortions.
4.1 Dataset and Feature Training